# Fast Decentralized State Estimation for Legged Robot Locomotion via EKF and MHE

### v1

@@ VI Evaluation
Now we evaluate the proposed decentralized estimation algorithm on our custom-designed hopping robot PogoX, the commercial quadrupedal robot Unitree Go1 and the bipedal robot Cassie. Both the EKF and MHE are implemented using C++ in ROS2 environment. The QP is solved using OSQP [27], and the VO is implemented via the open-source ORB-SLAM3 [18]. The MHEs are solved at 200 Hz, with a window size of 20 and 3-4 camera frames contained, without extensive effort on optimizing the implementation.
VI-A Hopping Robot PogoX
@@
We demonstrate our state estimation algorithm on the robot PogoX developed in [12], which is shown in Fig. 4 (a). We upgraded the mechanical design and the sensors of the original PogoX. This new version of PogoX weighs 3.65 kg, with a height of 0.53 m, and it has a quadcopter frame with a diagonal axle distance of 500 mm. The robot is equipped with a Garmin LIDAR-Lite v3HP LED distance measurement sensor, with an accuracy of ±5 cm and an update rate of 200 Hz. It is used for measuring the distance between the robot and the ground, equivalent to the leg kinematics on other legged robots. We use the Intel Realsense Depth Camera D455 for the onboard vision and an NGIMU with a gyroscope range of 2000 deg/s and an accelerometer sensitivity of 16 gs for inertial measurement. All state estimations are performed in real-time on an onboard Intel NUC 13 Pro (i7 13th Gen 1340P, P-core 5.00 GHz, E-core 3.70 GHz). Fig. 5 shows the block diagram of the estimation.
@@
For indoor testing, we use a motion-capture system to obtain the ground truth measurement of the velocity. Snapshots of experiments are presented in Fig. 4 (b). The robot is controlled to follow a square path in a clockwise direction. Fig. 6 presents the estimated robot velocity and orientation. The results show that our state estimation closely matches the ground truth. Compared with the widely used VIO OpenVINS [14], where the same noise covariances are used, the RMSEs of our method are reasonably better. For outdoor experiments, our robot can stably perform omnidirectional dynamic hopping on hard ground and soft grass, relying only on the onboard sensors and computation. This shows that our algorithm provides accurate and fast online state estimation.
@@
Figure 6: Online velocity and orientation estimation of PogoX experiment. The RMSEs of ours and OpenVINS for horizontal velocity are 0.1351 m/s and 0.1380 m/s. The RMSEs for orientation are 0.0424 rad and 0.0771 rad.
@@
VI-B Bipedal Cassie and Quadrupedal Unitree Go1
We evaluate our algorithms on Cassie and Go1. Our method shows better performance than the state-of-the-art estimation algorithms. We first validate our estimation on the simulation dataset [3] of Cassie. Fig. 7 shows the comparison result between our method, IEKF[3], and the ground truth. Using the same noise parameters, our method performs better than IEKF without incorporating Visual Information. We also tested our estimation algorithms on Go1, with the ground truth coming from the motion-capture system. During the experiment, we used the robot IMU along with Realsense Depth Camera D455. From Fig. 8, we can see that our method closely matches the ground truth by combining the visual information from VO, and achieves better performance compared with EKF fusion of IMU and Leg Odometry [2]. Moreover, with this decentralized framework, we find it significantly easier in practice to sequentially tune the orientation EKF and then the linear MHE than the centralized estimators such as the nonlinear EKF/IEFK.
Figure 8: Estimation results on Go1. The RMSEs of the linear velocities of our method and EKF are 0.0654 m/s and 0.1122 m/s, respectively.
@@ VII Conclusions and Future Work
In the future, we aim to further exploit the characteristic of QP that has a closed-form solution by utilizing numerical techniques to accelerate computation. We are also interested in exploiting the advantages of windowed optimization, to estimate robot states not limited to the floating base state. With the inclusion of physical constraints, the framework can be extended to the estimation of contact and joint torques for general robots during locomotion and manipulation tasks.

### v3

@@ VI Evaluation
Now we evaluate the proposed decentralized estimation algorithm on our custom-designed hopping robot PogoX, the commercial quadrupedal robot Unitree Go1 and the bipedal robot Cassie. Both the EKF and MHE are implemented using C++ in ROS2 environment. The MHE is solved using OSQP [28], and the VO is implemented via the open-source ORB-SLAM3 [19]. The MHEs are solved at 200 Hz, with a window size of 20 and 3-4 camera frames contained, without extensive effort on optimizing the implementation. The software implementation is open-sourced at [14]. For indoor testing, we use a motion-capture system consisting of 12 OptiTrack Prime-13s to obtain the ground truth at 200Hz. In legged robot estimation, yaw is typically unobservable, so velocity estimates in the following sections are given in the body frame rather than the world frame.
@@
Hopping Robot PogoX: We demonstrate our state estimation algorithm on the robot PogoX developed in [12], which is shown in Fig. 4 (a). We upgraded the mechanical design and the sensors of the original PogoX. This new version of PogoX weighs 3.65 kg, with a height of 0.53 m, and it has a quadcopter frame with a diagonal axle distance of 500 mm. The robot is equipped with a Garmin LIDAR-Lite v3HP LED distance measurement sensor, with an accuracy of ±5 cm and an update rate of 200 Hz. It is used for measuring the distance between the robot and the ground, equivalent to the leg kinematics on other legged robots. We use the Intel Realsense Depth Camera D455 for the onboard vision and an NGIMU with a gyroscope range of 2000 deg/s and an accelerometer sensitivity of 16 gs for inertial measurement. All state estimations are performed in real-time on an onboard Intel NUC 13 Pro (i7 13th Gen 1340P, P-core 5.00 GHz, E-core 3.70 GHz). Fig. 5 shows the block diagram of the estimation.
@@
Snapshots of indoor experiments are presented in Fig. 4 (b). The robot is controlled to follow a square path in a clockwise direction. Fig. 6 presents the estimated robot velocity and orientation. The results show that our state estimation closely matches the ground truth. Compared with the widely used VIO OpenVINS [15], where the same noise covariances are used, the RMSEs of our method are reasonably better. Note that the distance LIDAR does not contribute sufficiently to the velocity estimation due to its low precision measurements, yet it is necessary to include to estimate the vertical height that is used for hopping control. Including the distance LIDAR reduced the RMSE of height estimation from 1.1095m to 0.0329m. The brief ground truth outages in Fig. 6 are caused by marker occlusion. For outdoor experiments, our robot can stably perform omnidirectional dynamic hopping on hard ground and soft grass, relying only on the onboard sensors and computation. This shows that our algorithm provides accurate and fast online state estimation.
Bipedal Cassie and Quadrupedal Unitree Go1: We evaluate our algorithms on Cassie and Go1. Our method shows better performance than the state-of-the-art estimation algorithms. We first validate our estimation on the simulation dataset [4] of Cassie. Fig. 7 shows the comparison result between our method, IEKF[4], and the ground truth. Using the same sensors and noise parameters, our method performs better than IEKF. We also test our estimation algorithms on Go1. During the experiment, we attach an additional Realsense camera D455 to the torso. Fig. 8 and Table. I show that our method closely matches the ground truth, and achieves better performance compared with EKF [2] and IEKF [4] that fuse IMU and Leg Odometry.
@@
Figure 6: Online velocity and orientation estimation of PogoX experiment. The RMSEs of ours and OpenVINS for horizontal velocity are 0.1351 m/s and 0.1380 m/s. The RMSEs for orientation are 0.0424 rad and 0.0771 rad. Here the LO implies the distance LIDAR.
@@
MHE Window Size Analysis: We investigate the MHE error and computation time in Fig. 9 by increasing the window size from 1 to 20, on both Go1 and PogoX. The RMSE improves with larger windows due to the inclusion of more VO frames; yet errors can rise as VO measurements are discontinuous, especially around the window size of 5. As window size increases beyond 10, the RMSE becomes insensitive to the increase in windows, as sufficient VO frames have been included already. This result highlights the value of MHE in using windowed measurements for precise estimation.
Table I: RMSEs of the estimations on Go1 hardware.
Estimation Our Method Our Method EFK IEKF
Method (w/o VO)
$\text{RMSE}_{v}$ [m/s] 0.0654 0.0847 0.0901 0.0817
$\text{RMSE}_{\text{Euler}}$ [rad] 0.0160 0.0230 0.0218 0.0157
Figure 8: Estimation results on Unitree Go1 hardware.
Figure 9: Window size effects on (a) estimation RMSE of linear velocities and (b) computation cost.
@@ VII Conclusions and Future Work
In the future, we aim to exploit the characteristics of QP that have a closed-form solution by utilizing numerical techniques to accelerate computation. We will also conduct theoretical and experimental comparisons on arrival cost calculation between our method and other existing approaches. Additionally, we are interested in exploring different orientation estimators and leveraging windowed optimization to estimate robot states beyond the floating base state. By incorporating physical constraints, we seek to extend the framework to include contact and joint torque estimation in general robots during locomotion and manipulation tasks.